HUBBLE SPACE TELESCOPE DAILY REPORT #1455 PERIOD COVERED: 0000Z (UTC) 08/15/95 - 0000Z (UTC) 08/16/95 Daily Status Report as of 228/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed FGS 6158 (Parallaxes of Hyades Cluster Members, Cycle 5 - Part 2) The FGSs were used to make astrometric POS Mode observations (Part Two) of members (set PCEN645, VA645, AST121-REF, et.al.) of the Hyades Cluster in order to measure their parallaxes. This is part two of proposal #6157. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed Two Sets of WF/PC-2 6189 (WF/PC-2 Cycle 5 Calibration: Internal Monitor - Vis Flats) The WF/PC-2 was used to make internal visual flat calibration images. Both photometric and narrow-band filters were used. The observations were executed as planned, and no problems were reported. 1.3 Completed GHRS/2 6327 (Carbon Monoxide on Mars) The High Resolution Spectrograph (Side Two) was used to make spectrographic observations of the planet Mars to study the carbon monoxide content (i.e., the Cameron Band emissions) of its atmosphere. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed WF/PC-2 5932 (Morphologies of High Redshift Radio Galaxies) The WF/PC-2 was used to make observations the high redshift galaxy GAL-1413-0 ( At distance Z>2) in order to study its morphology, especially the fine detail in clumping. The images were obtained in the continuum and Lyman-Alpha. The observations were executed as planned, and no anomalies were noted. 1.5 Completed Two Sets of FOS/RD 6236 (FOS Cycle 5: Location of Spectra and Wavelength Calibration: Red Side Detector) The Faint Object Spectrograph (Red Side) was used to make calibration observations in order to determine the location of the spectra and to perform a wavelength calibration. All observations were internal. A Taled calibration was also completed. The observations were completed as planned, and no anomalies were reported. 1.6 Completed WF/PC-2 6190 (WF/PC-2 Cycle 4 Calibration: Internal Flats) The WF/PC-2 was used to make an internal flat filter sweep and an internal preflash level test. The internal lamp is also used as a source for the preflash test. The observation was executed as scheduled, and no anomalies were noted. 1.7 Completed WF/PC-2 6250 (WF/PC-2 Cycle 5 Calibration: Internal Monitor) The WF/PC-2 was used to make internal bias, internal flat images and K-Spot calibrations to monitor the health of the cameras. The observations were executed as scheduled, and no problems were reported. 1.8 Completed FOS/RD 5825 (The Structure of Triton's Lower Atmosphere) The Faint Object Spectrograph (Red Side) was used to make Rapid Mode spectrographic observations of Neptune's moon Triton to study its lower atmosphere during an occultation with a 12.9 magnitude star. The observations were completed as planned, and no problems were noted. 1.9 Completed Three Sets of WF/PC-2 6253 (Pure Parallel Archive Program) The WF/PC-2 was used to execute a pure parallel program taking ultraviolet observations running concurrently with prime science observations. The intent is to search for extra-galactic sources with flux excesses at 3000 A0 and galactic sources with excesses at 1600 A0. The observations were completed as planned, and no problems were reported. 1.10 Completed GHRS/1/2 5879 (The Deuterium to Hydrogen Ration in the Interstellar Medium) The High Resolution Spectrograph (Sides One and Two) was used to make spectrographic observations of the late type star HD4128 in order to study the interstellar medium. The objective is to determine the deuterium to hydrogen ratio in the interstellar medium in the line of sight to this star. At 227/191153Z, GHRS status buffer messages 20 (number of correction slews exceeded in target locate) and 15 (coarse x-null balance failure) occurred (HSTAR #5251). While OPUS feels the messages may have been caused by a diffuse target, any possible effects on science are not currently known. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 10 Successful : 10 Scheduled Reacquisitions : 5 Successful : 5 2.2 FHST Updates: Scheduled : 18 Successful : 18 2.3 Operations Notes: A series of ground system problems occurred during the night. At 227/1021Z, the PRS #1 and TAC #1 systems crashed. At 228/0655Z, the PRS backup system also crashed. DOC personnel were unable to restore either the prime or backup systems, so operations were transferred to the SAMS facility at 228/0950Z. Both Code 500 and LORAL personnel have been notified. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-